Spartan Credit Bank is offering 7.4 percent compounded daily on its savings accounts. You deposit $5,800 today. a. How much will you have in the account in 6 years? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) b. How much will you have in the account in 11 years? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) c. How much will you have in the account in 18 years?
Answer #1

a) Amount in 6 years:

=FV(rate,nper,pmt,pv)

=FV(7.4%/365,6*365,,-5800)

=9041.39

b) Amount in 11 years:

=FV(7.4%/365,11*365,,-5800)

=13089.04

c) Amount in 18 years:

=FV(7.4%/365,18*365,,-5800)

=21970.99